# Cohomology of modular form connections on complex curves

### v3

@@ 2.1. Chain complex spaces of $n$-variable modular forms
Modular functions depending on $n$ variables of the space $\left\{\mathcal{Z}\left({\bf z}_{n},\mu\right),n\geq 0\right\}$ form a space of connections depending on $n$ points. For $n\geq 0$, the reduction cohomology of a compact complex curve of genus $g$ is $H^{n}(\mu)={\mathcal{C}on}^{n}/G^{n-1}$.
@@
Thus, the formula (2.2) gives (2.3). ∎
The geometrical meaning of (2.5) consists in the following. Due to modular properties of functions $\mathcal{Z}({\bf z}_{n},\mu)$ depending on $n$ variables, (2.5) is also interpreted as relations among modular forms. The condition (2.3) defines a complex variety in ${\bf z}_{n}\in\mathbb{C}^{n}$. As most identities (e.g., trisecant identity [12, 35] and triple product identity [31]) for modular functions depending on $n$ variables (2.5) has its algebraic-geometrical meaning. The condition (2.5) relates finite series of modular functions on $M$ with rational function coefficients (at genus $g=0$) [50], or (deformed) elliptic functions (at genus $g=1$) [50, 31], or generalizations of classical elliptic functions (at genus $g\geq 2$) [20, 45].
2.3. Cohomology
In this Section we compute the reduction cohomology defined by (2.6)–(2.7). The main result of this paper is the following
@@ 2.3. Cohomology
The $n$-th reduction cohomology of the spaces $C^{n}(\mu)$ (2.1) of modular forms $\mathcal{Z}({\bf z}_{n},\mu)$ is the space of recursively generated (by the reduction formulas (2.3)) functions with $z_{i}\notin{\mathfrak{Z}}_{i}$, for $1\leq i\leq n$, satisfying the condition
@@
The first cohomology is given by the space of transversal (i.e., with vanishing sum over $q$, $q^{\prime}$) one-point connections $\mathcal{Z}\left(x_{1},\mu\right)$ provided by coefficients in terms of series of special functions. The second cohomology is given by a space of generalized higher genus complex kernels corresponding to $M$.
@@
By definition (2.7), the $n$-th reduction cohomology is defined by the subspace of $C^{n}(\mu)$ of functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ satisfying (2.9) modulo the subspace of $C^{n}(\mu)$ modular functions depending on $n$ variables $\mathcal{Z}\left({\bf z}^{\prime}_{n},\mu\right)$ resulting from:
@@
We assume that, subject to other fixed $\mu$-parameters, modular functions depending on $n$ variables are completely determined by all choices ${\bf z}_{n}\in\mathbb{C}^{n}$. Thus, the reduction cohomology can be treated as depending on set of ${\bf z}_{n}$ only with appropriate action of endomorphisms generated by $z_{n+1}$. Consider a non-vanishing solution $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ to (2.9) for some ${\bf z}_{n}$. Let us use the reduction formulas (2.3) recursively for each $z_{i}$, $1\leq i\leq n$ of ${\bf z}_{n}$ in order to express $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ in terms of modular form without parameters $\mathcal{Z}\left(\mu\right)$, i.e., we obtain
$\mathcal{Z}\left({\bf z}_{n},\mu\right)={\mathcal{D}}({\bf z}_{n},\mu)\mathcal{Z}\left(\mu\right),$ (2.11)
as in [31]. It is clear that $z_{i}\notin{\mathfrak{Z}}_{i}$ for $1\leq i\leq n$, i.e., at each stage of the recursion procedure towards (2.11), otherwise $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ would be zero. Thus, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ is explicitly known and is repsented by a series of auxiliary functions ${\mathcal{D}}({\bf z}_{n})$ depending on moduli space parameters $\mu$. Consider now $\mathcal{Z}\left({\bf z}^{\prime}_{n}\right)$ given by (2.10). It is either vanishes when $z_{n-i}\in{\mathfrak{Z}}_{n-i}$, $2\leq i\leq n$, or given by (2.11) with ${\bf z}^{\prime}_{n}$ arguments. The general idea of deriving reduction formulas is to consider the double integration of $\mathcal{Z}\left({\bf z}_{n}\right)$ along small circles around two auxiliary variables with the action of reproduction kernels inserted. Then, these procedure leads to recursion formulas relating $\mathcal{Z}({\bf z}_{n+1},\mu)$ and $\mathcal{Z}({\bf z}_{n},\mu)$ with functional coefficients depending on the nature of corresponding modular functions, and $M$. In [48, 31] formulas to modular functions depending on $n$ variables were explicitly and recursively obtained in various specific examples. In terms of $z_{n+1}$, we are able to transfer in (2.9) the action of $T_{l,k,m}(\mu)$-operators into an analytical continuation of $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ multi-valued holomorphic functions to domains $D_{n}\subset M$ with $z_{i}\neq z_{j}$ for $i\neq j$. Namely, in (2.9), the operators $T_{l,k,m}(\mu)$ shift the formal parameters ${\bf z}_{n}$ by $z_{n+1}$, i.e., ${\bf z}^{\prime}_{n}={\bf z}_{n}+z_{n+1}$. Thus, the $n$-th reduction cohomology is given by the space of analytical continuations of modular functions depending on $n$ variables $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ with ${\bf z}_{n-1}\notin{\mathfrak{Z}}_{n-1}$ that are solutions of (2.9). ∎
@@ 3. Motivating examples
In this Section we provide the examples of sets of coefficients $f_{l,k,m}$, operators $T_{l,k,m}(\mu)$, and functions $\mathcal{Z}({\bf z}_{n},\mu)$ satisfying the reduction formulas. Functions $\mathcal{Z}({\bf z}_{n},\mu)$ satisfy certain automorphic properties with respect to corresponding groups [50, 31, 20, 45]. It is easy to see that all reduction formulas of this Section are in the form (2.2). These examples motivate the general construciton of the complex in Section 2. The reduction cohomology depends on the kind of modular forms (via moduli parameters which we denote $\mu$) and genus of $M$. As it was shown in [33, 26, 27], existence of reduction formulas is related in some sense to modularity.
@@ 3.1. The rational case
In (cf., e.g., [50]) we find for the rational case functions depending on $n$ variables, the reduction formulas
@@
where $\iota_{z,w}:\mathbb{C}[z_{1},\ldots,z_{n}]\to\mathbb{C}[[z_{1},z_{1}^{-1}\ldots,z_{n}z_{n}^{-1}]]$ are certain maps [13]. Let us take $z_{n+1}$ as the variable of expansion. Then the $n$-th reduction cohomology $H^{n}(\mu)$ is given by the space of rational functions recursively generated by (2.3) with ${\bf z}_{n}\notin{\mathfrak{Z}}_{n}$, satisfying (2.9) with rational function coefficients $f_{l,k,m}(z_{n+1},z_{k})$, and modulo the space of functions depending on $n$ variables obtained by the recursion procedure, not given by $\delta^{n-1}\mathcal{Z}({\bf z}_{n-1},\mu)$. It is possible to rewrite (2.9), in the form
@@
which is an equation for an analytical continuation of $\mathcal{Z}({\bf z}_{n}+(z_{n+1})_{k},\mu)$ with different functions $\widetilde{f}_{k,m}$. Using the reduction formulas (2.3) we obtain $\mathcal{Z}({\bf z}_{n}+(z_{n+1})_{k},\mu)={\mathcal{D}}({\bf z}_{n+1},\mu)$, where ${\mathcal{D}}({\bf z}_{n+1},\mu)$ is given by the series of rational-valued functions in ${\bf z}_{n+1}\notin{\mathfrak{Z}}_{n}$ resulting from the recursive procedure starting from function depending on $n$ variables to the partition function. Thus, in this example, the $n$-th cohomology is the space of analytic extensions of rational function solutions to the equation (2.9) with rational function coefficients.
@@ 3.2. Modular and elliptic functions
$P_{m+1,\lambda}\left(w,\tau\right)$ $=$ $\frac{(-1)^{m+1}}{m!}\sum_{n\in\mathbb{Z}\backslash\{-\lambda\}}\frac{n^{m}q_{w}^{n}}{1-q^{n+\lambda}}.$ (3.2)
@@ 3.4. The case of deformed elliptic functions
where $p(n,k)$ is a parity multiplier. A deformed Weierstrass function is defined as follows [10, 31]. Let $(\theta,\phi)\in U(1)\times U(1)$ denote a pair of modulus one complex parameters with $\phi=\exp(2\pi i\lambda)$ for $10\leq\lambda<$ 1. For $z\in\mathbb{C}$ and $\tau\in\mathbb{H}$ we define deformed Weierstrass functions for $k\geq 1$,
@@
3.5. The reduction formulas for Jacobi functions
In this Subsection we recall the reduction formulas derived in [31, 7]. For $\alpha\in\mathbb{C}$, we now provide the following reduction formula for formal Jacobi functions depending on $n$ variables. Let ${\bf z}_{n+1}\in\mathbb{C}^{n+1}$, $\alpha\in\mathbb{C}$. For $\alpha z\notin{\mathbb{Z}\tau}+\mathbb{Z}$, we have
@@ 3.5. The reduction formulas for Jacobi functions
with $P_{m+1,\lambda}\left(w,\tau\right)$ defined in (3.2). Next we provide the reduction formula for Jacobi functions depending on $n$ variables. For $l\geq 1$ and $\alpha z\notin{\mathbb{Z}\tau}+\mathbb{Z}$, we have
@@ 3.6. Multiparameter Jacobi forms
For multiparameter Jacobi forms [11, 49, 26, 27, 7], the reduction formulas are found using an analysis that is similar to that in [50, 31]. The following two formulas reduce any multiparameter Jacobi function depending on $n$ variables to a linear combination of Jacobi functions depending on $n-1$ variables with modular coefficients. For each $1\leq j\leq m$ we have
$\mathcal{Z}({\bf z}_{n+1},\mu,\tau)=\delta_{{\bf z}_{n}\cdot(\alpha)_{n},\mathbb{Z}}\;\mathcal{Z}({\bf z}_{n},(\alpha)_{n},\mu(m))+\sum_{s=1}^{n}\sum_{k\geq 0}\tilde{P}_{k+1}(z_{s}-z,{\bf z}_{n}\cdot(\alpha_{n}),\tau)\mathcal{Z}({\bf z}_{n},\mu_{s,k}\tau),$
where $\delta_{{\bf z}\cdot(\mu)_{n},\mathbb{Z}}$ is $1$ if ${\bf z}_{n}\cdot(\mu)_{n}\in\mathbb{Z}$ and is 00 otherwise. With the same assumptions as above, for $p\geq 1$,
@@ 3.7. The genus two counterparts of Weierstrass functions
In this Subsection we recall the definition of genus two Weierstrass functions [20]. For $m$, $n\geq 1$, we first define a number of infinite matrices and row and column vectors: $\Gamma(m,n)=\delta_{m,-n+2p-2}$, $\Delta(m,n)=\delta_{m,n+2p-2}$. We also define the projection matrix $\Pi=\Gamma^{2}=\begin{bmatrix}\mathbbm{1}_{2p-1}&0\\ 0&\ddots\end{bmatrix}$, where ${\rm Id}_{2p-3}$ denotes the $32p-$ 3 dimensional identity matrix and ${\rm Id}_{-1}=0$. Let $\Lambda_{a}$ for $a\in\{1,2\}$ be the matrix with components
@@
Introduce also an infinite row vector $\mathbb{Q}(p;x)=\mathbb{R}(x)\Delta\left(\mathbbm{1}-\widetilde{\Lambda}_{\overline{a}}\widetilde{\Lambda}_{a}\right)^{-1}$, for $x$ on the torus. Notice that $\widetilde{\Lambda}_{a}=\Lambda_{a}\Delta$. One introduces $\mathbb{P}_{j+1}(x)$ $=$ $\frac{(-1)^{j}}{j!}$ $\mathbb{P}_{1}(x)$, and $j\geq 0$, is the column with components
@@ 3.8. The genus two case
In this Subsection we recall [20] the construction and reduction formulas for modular functions defined on genus two complex curve. In particular, we use the geometric construction developed in [48]. For a complex parameter $\epsilon=z_{1}z_{2}$, the modular form with no dependence on parameters is defined on a genus two complex curve by
@@
where parameters $\mu_{1}$ and $\mu_{2}$ are related. We then recall [20] the formal genus two reduction formulas for modular functions depending on $n$ variables. Let $x_{n+1}$, ${\bf y}_{k}$ and ${\bf y}^{\prime}_{l}$ be inserted on two torai. We consider the genus two modular function depending on $n$ variables $\mathcal{Z}\left(z_{n+1},{\bf z}_{k};{\bf z}^{\prime}_{l},\mu)\right)=\sum_{r\geq 0}\epsilon^{r}\mathcal{Z}\left(z_{n+1},{\bf x}_{k},\mu_{1},\tau_{1}\right)\;\mathcal{Z}\left({\bf x}^{\prime}_{l},\mu_{2},\tau_{2}\right)$, where the sum as in (3.10). First, one defines the functions $\mathcal{Z}_{n,a}$ for $a\in\{1,2\}$, via elliptic quasi-modular forms
@@
In [20] it is proven that the genus two function depending on $n=k+l$ variables inserted at $x_{n-k}$, ${\bf y}_{k}$ on two torai has the following reduction formula

### v5

@@ 2.2. Recursion relations as mutipoint connections
Correlation functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$, $n\geq 0$, belong to the space of multipoint connections. The corresponding recursion cohomology is given by $H^{n}(\mu)={\mathcal{C}on}^{n}/G^{n-1}$.
@@
Thus, the formula (2.2) gives (LABEL:reduction). ∎
Now let us describe the geometric meaning of the relations (2.5). In conformal field theory, identities coming from computations, such as the triple product identity [32] and the trisecant identity [13, 36], arrise from computations of correlation functions via recursion formulas. In many known examples correlation functions $\mathcal{Z}({\bf z}_{n},\mu)$ enjoy modular properties. Therefore, (2.5) is also interpreted as relations among modular forms. The condition (LABEL:reduction) defines a complex variety in ${\bf z}_{n}\in\mathbb{C}^{n}$. The conditions (2.5) make a connection among finite series of correlation functions on $M$ with rational function coefficients (at genus $g=0$) [53], of deformed elliptic functions (at genus $g=1$) [53, 32], or generalizations of classical elliptic functions (at genus $g\geq 2$) [21, 46].
2.3. Computation of the recursion cohomology
The main result of this paper is
@@ 2.3. Computation of the recursion cohomology
The $n$-th recursion cohomology of the spaces $C^{n}(\mu)$ (LABEL:cnmu) of correlation functions $\mathcal{Z}({\bf z}_{n},\mu)$ is the space of functions recursively generated by (LABEL:reduction) with $z_{i}\notin{\mathfrak{Z}}_{i}$, for $1\leq i\leq n$, and satisfying the condition
@@
Note that the first cohomology is given by the space of transversal one-point connections $\mathcal{Z}\left(x_{1},\mu\right)$ with coefficients in terms of special functions. The spaces of higher genus generalized complex reproduction kernels corresponding to $M$ represent the second cohomology.
@@
According definition (LABEL:pusto), the $n$-th recursion cohomology is a coset of the subspace correlation functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ satisfying (2.9) with respect to a subspace of $C^{n}(\mu)$ of correlation functions $\mathcal{Z}\left({\bf z}^{\prime}_{n},\mu\right)$ resulting from:
@@
With other $\mu$-parameters fixed, correlation functions are completely determined by all choices ${\bf z}_{n}\in\mathbb{C}^{n}$. The recursion cohomology is a function of ${\bf z}_{n}$ only with appropriate recursion formulas adition of the parameter $z_{n+1}$. For a non-zero solution $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ for (2.9) with some ${\bf z}_{n}$, apply the recursion formulas (LABEL:reduction) recursively for each $z_{i}$, $1\leq i\leq n$ of ${\bf z}_{n}$. Thus, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ is then given as an expression in the correlation function without parameters $\mathcal{Z}\left(\mu\right)$. We have, similar to [32],
$\mathcal{Z}\left({\bf z}_{n},\mu\right)={\mathcal{D}}({\bf z}_{n},\mu)\mathcal{Z}\left(\mu\right).$ (2.11)
For $z_{i}\notin{\mathfrak{Z}}_{i}$, $1\leq i\leq n$, i.e., at each step of the recursion leading to (LABEL:topaz), or, otherwise $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ vanishes. As a result, $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ we obtain an explicitly formula a series of functions ${\mathcal{D}}({\bf z}_{n})$ of the moduli space parameters $\mu$. The function $\mathcal{Z}\left({\bf z}^{\prime}_{n}\right)$ given by (2.10) vanishes for $z_{n-i}\in{\mathfrak{Z}}_{n-i}$, $2\leq i\leq n$, or is commes from (LABEL:topaz) with parameters ${\bf z}^{\prime}_{n}$ . Recursion formulas relating the correlation functions $\mathcal{Z}({\bf z}_{n+1},\mu)$ and $\mathcal{Z}({\bf z}_{n},\mu)$ with functional coefficients were obtained as a result of the double integration of the action of reproduction kernels on $\mathcal{Z}\left({\bf z}_{n}\right)$ via small paths around two auxiliary points. We can move now the action of $T_{l,k,m}(\mu)$-operators in (2.9) to an analytical continuation of $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ which are multivalued holomorphic functions on domains $D_{n}\subset M$ with $z_{i}\neq z_{j}$, $i\neq j$. In (2.9) the operators $T_{l,k,m}(\mu)$ shift the formal parameters ${\bf z}_{n}$ by $z_{n+1}$, i.e., ${\bf z}^{\prime}_{n}={\bf z}_{n}+z_{n+1}$. As an upshot, the space of analytical continuations of correlation functions $\mathcal{Z}\left({\bf z}_{n},\mu\right)$ with ${\bf z}_{n-1}\notin{\mathfrak{Z}}_{n-1}$ which are solutions of (2.9) gives the $n$-th recursion cohomology. ∎
@@ 3. Motivating examples
In this Section we provide the examples of sets of coefficients $f_{l,k,m}$, operators $T_{l,k,m}(\mu)$, and functions $\mathcal{Z}({\bf z}_{n},\mu)$ satisfying the recursion formulas. Functions $\mathcal{Z}({\bf z}_{n},\mu)$ satisfy certain automorphic properties with respect to corresponding groups [53, 32, 21, 46]. It is easy to see that all recursion formulas of this Section are in the form (LABEL:poros). These examples motivate the general construciton of the complex in Section ??. The recursion cohomology depends on the kind of correlation functions (via moduli parameters which we denote $\mu$) and genus of $M$. As it was shown in [34, 27, 28], existence of recursion formulas is related in some sense to modularity.
@@ 3.1. The rational case
In (cf., e.g., [53]) we find for the rational case functions depending on $n$ variables, the recursion formulas
@@
where $\iota_{z,w}:\mathbb{C}[z_{1},\ldots,z_{n}]\to\mathbb{C}[[z_{1},z_{1}^{-1}\ldots,z_{n}z_{n}^{-1}]]$ are certain maps [14]. Let us take $z_{n+1}$ as the variable of expansion. Then the $n$-th recursion cohomology $H^{n}(\mu)$ is given by the space of rational functions recursively generated by (LABEL:reduction) with ${\bf z}_{n}\notin{\mathfrak{Z}}_{n}$, satisfying (2.9) with rational function coefficients $f_{l,k,m}(z_{n+1},z_{k})$, and modulo the space of functions depending on $n$ variables obtained by the recursion procedure, not given by $\delta^{n-1}\mathcal{Z}({\bf z}_{n-1},\mu)$. It is possible to rewrite (2.9), in the form
@@
which is an equation for an analytical continuation of $\mathcal{Z}({\bf z}_{n}+(z_{n+1})_{k},\mu)$ with different functions $\widetilde{f}_{k,m}$. Using the recursion formulas (LABEL:reduction) we obtain $\mathcal{Z}({\bf z}_{n}+(z_{n+1})_{k},\mu)={\mathcal{D}}({\bf z}_{n+1},\mu)$, where ${\mathcal{D}}({\bf z}_{n+1},\mu)$ is given by the series of rational-valued functions in ${\bf z}_{n+1}\notin{\mathfrak{Z}}_{n}$ resulting from the recursive procedure starting from function depending on $n$ variables to the partition function. Thus, in this example, the $n$-th cohomology is the space of analytic extensions of rational function solutions to the equation (2.9) with rational function coefficients.
@@ 3.2. Modular and elliptic functions
[…]
@@ 3.4. The case of deformed elliptic functions
where $p(n,k)$ is a parity multiplier. A deformed Weierstrass function is defined as follows [11, 32]. Let $(\theta,\phi)\in U(1)\times U(1)$ denote a pair of modulus one complex parameters with $\phi=\exp(2\pi i\lambda)$ for $0\leq\lambda<1$. For $z\in\mathbb{C}$ and $\tau\in\mathbb{H}$ we define deformed Weierstrass functions for $k\geq 1$,
@@
3.5. The recursion formulas for Jacobi functions
In this Subsection we recall the recursion formulas derived in [32, 8]. For $\alpha\in\mathbb{C}$, we now provide the following recursion formula for formal Jacobi functions depending on $n$ variables. Let ${\bf z}_{n+1}\in\mathbb{C}^{n+1}$, $\alpha\in\mathbb{C}$. For $\alpha z\notin{\mathbb{Z}\tau}+\mathbb{Z}$, we have
@@ 3.5. The recursion formulas for Jacobi functions
with $P_{m+1,\lambda}\left(w,\tau\right)$ defined in (3.2). Next we provide the recursion formula for Jacobi functions depending on $n$ variables. For $l\geq 1$ and $\alpha z\notin{\mathbb{Z}\tau}+\mathbb{Z}$, we have
@@ 3.6. Multiparameter Jacobi forms
For multiparameter Jacobi forms [12, 52, 27, 28, 8], the recursion formulas are found using an analysis that is similar to that in [53, 32]. The following two formulas reduce any multiparameter Jacobi function depending on $n$ variables to a linear combination of Jacobi functions depending on n1n-1 variables with modular coefficients. For each $1\leq j\leq m$ we have
$\mathcal{Z}({\bf z}_{n+1},\mu,\tau)$ $=$ $\delta_{{\bf z}_{n}\cdot(\alpha)_{n},\mathbb{Z}}\;\mathcal{Z}({\bf z}_{n},(\alpha)_{n},\mu(m))$
$+$ $\sum_{s=1}^{n}\sum_{k\geq 0}\tilde{P}_{k+1}(z_{s}-z,{\bf z}_{n}\cdot(\alpha_{n}),\tau)\mathcal{Z}({\bf z}_{n},\mu_{s,k}\tau),$
where $\delta_{{\bf z}\cdot(\mu)_{n},\mathbb{Z}}$ is $1$ if ${\bf z}_{n}\cdot(\mu)_{n}\in\mathbb{Z}$ and is 0 otherwise. With the same assumptions as above, for $p\geq 1$,
@@ 3.7. The genus two counterparts of Weierstrass functions
In this Subsection we recall the definition of genus two Weierstrass functions [21]. For $m$, $n\geq 1$, we first define a number of infinite matrices and row and column vectors: $\Gamma(m,n)=\delta_{m,-n+2p-2}$, $\Delta(m,n)=\delta_{m,n+2p-2}$. We also define the projection matrix $\Pi=\Gamma^{2}=\begin{bmatrix}\mathbbm{1}_{2p-1}&0\\ 0&\ddots\end{bmatrix}$, where ${\rm Id}_{2p-3}$ denotes the 2p32p-3 dimensional identity matrix and ${\rm Id}_{-1}=0$. Let $\Lambda_{a}$ for $a\in\{1,2\}$ be the matrix with components
@@
Introduce also an infinite row vector $\mathbb{Q}(p;x)=\mathbb{R}(x)\Delta\left(\mathbbm{1}-\widetilde{\Lambda}_{\overline{a}}\widetilde{\Lambda}_{a}\right)^{-1}$, for $x$ on the torus. Notice that $\widetilde{\Lambda}_{a}=\Lambda_{a}\Delta$. One introduces $\mathbb{P}_{j+1}(x)$ == $\frac{(-1)^{j}}{j!}$ $\mathbb{P}_{1}(x)$, and $j\geq 0$, is the column with components
@@ 3.8. The genus two case
In this Subsection we recall [21] the construction and recursion formulas for correlation functions defined on genus two complex curve. In particular, we use the geometric construction developed in [51]. For a complex parameter $\epsilon=z_{1}z_{2}$, the correlation function with no dependence on parameters is defined on a genus two complex curve by
@@
where parameters $\mu_{1}$ and $\mu_{2}$ are related. We then recall [21] the formal genus two recursion formulas for correlation functions depending on $n$ variables. Let $x_{n+1}$, ${\bf y}_{k}$ and ${\bf y}^{\prime}_{l}$ be inserted on two torai. We consider the genus two correlation function depending on $n$ variables $\mathcal{Z}\left(z_{n+1},{\bf z}_{k};{\bf z}^{\prime}_{l},\mu)\right)=\sum_{r\geq 0}\epsilon^{r}\mathcal{Z}\left(z_{n+1},{\bf x}_{k},\mu_{1},\tau_{1}\right)\;\mathcal{Z}\left({\bf x}^{\prime}_{l},\mu_{2},\tau_{2}\right)$, where the sum as in (LABEL:eq:Zdef). First, one defines the functions $\mathcal{Z}_{n,a}$ for $a\in\{1,2\}$, via elliptic quasi-modular forms
@@
In [21] it is proven that the genus two function depending on $n=k+l$ variables inserted at $x_{n-k}$, ${\bf y}_{k}$ on two torai has the following recursion formula